# Floorplan-SLAM: A Real-Time, High-Accuracy, and Long-Term Multi-Session Point-Plane SLAM for Efficient Floorplan Reconstruction

### v1

@@ I INTRODUCTION
Indoor robot navigation and high-level scene understanding have become critical components in a variety of applications, ranging from service robots in assisted living environments to autonomous drones operating in complex indoor facilities. In these scenarios, generating a floorplan in real-time is particularly advantageous, as it not only provides geometric structure for accurate localization but also furnishes semantic cues for higher-level decision-making.
Figure 1: The real-time floorplan reconstruction results of Floorplan-SLAM on the meeting room sequence from the self-collected dataset. The scene captured over a two-day period consists of low-texture corridors and multiple rooms with many clutter objects.
Despite the growing need for real-time floorplan reconstruction, existing approaches often present significant limitations. Some methods operate in an offline manner, requiring a complete map of the environment prior to reconstruction [1, 2, 3]. Others rely on expensive sensors such as LiDAR or RGB-D cameras [4, 5], or computationally intensive neural networks [6, 7, 8], which increase both deployment costs and system complexity.
While monocular or stereo camera-based SLAM has attempted to use high-level features such as planes for real-time vectorized map construction [9, 10, 11], these methods lack high-level scene layout understanding, thereby failing to achieve complete and compact indoor vectorized reconstruction. Additionally, they cannot perform robust plane extraction in weakly textured scenes.
@@
In this context, an effective real-time solution that works with more accessible and cost-effective sensors, such as stereo cameras, remains uncertain. In this work, we present Floorplan-SLAM, a novel framework that tightly integrates floorplan reconstruction into a SLAM system by seamlessly interacting with plane extraction, pose estimation, and back-end optimization. Unlike prior methods, our approach operates exclusively with stereo cameras under standard computing resources, allowing real-time performance in large-scale indoor settings. A robust plane extraction algorithm is first developed to handle weakly textured surfaces by leveraging two spatially complementary features in a compact plane parameter space. We further introduce a vectorized reconstruction module tightly integrated with the SLAM system that formulates and solves an innovative optimization problem using the continuously optimized plane landmarks and poses from the SLAM system, enabling real-time incremental floorplan reconstruction. Furthermore, by leveraging the map merging capability of multi-session SLAM, our method supports long-term floorplan reconstruction across multiple sessions without redundant data collection, making it more efficient for large-scale environments.
@@
    We introduce a floorplan reconstruction module that is tightly coupled with the multi-session SLAM system, leveraging continuously optimized plane landmarks and poses from the SLAM system to formulate and solve a binary linear programming problem with novel trajectory constraint thereby enabling real-time, high-accuracy, and long-term floorplan reconstruction.
@@ II Related Work
With the development of neural networks, deep learning-based vectorized reconstruction methods have become of interest. These methods [6, 7, 8] typically first project the input dense point cloud onto a top view to create a 2D density map, and then use neural networks to infer the floorplan of the scene from the projected density map. However, these methods are only applicable to specific scenarios, exhibit limited generalizability, and incur high time cost and computational resource consumption.
To achieve real-time vectorized reconstruction, some works [9, 4, 5] utilized visual SLAM to extract plane features from image sequences, which are then used to create landmarks for map construction. However, these methods either lack high-level scene layout understanding [9], thereby failing to achieve complete and compact indoor vectorized reconstruction, or integrate the acquired poses and scene structures in a loosely coupled manner [4, 5], leading to poor system robustness due to the inability to perform mutual optimization between pose and scene structure. In addition, these methods can only process one set of consecutive images at a time, making it impossible to merge multiple sequences and achieve long-term mapping.
@@
Figure 2: System Overview. The proposed system consists of two stages: Stereo Point-Plane SLAM and Floorplan Reconstruction. The latter utilizes optimized plane landmarks and trajectory information to generate an accurate and well-structured floorplan.
@@ III Point-Plane-Based Stereo SLAM System
To transform and integrate the planar structure obtained from each frame into the global map, the camera pose of each frame needs to be estimated. Therefore, we propose a point-plane-based stereo SLAM system to estimate camera poses, while creating and managing plane landmarks, which are further processed for floorplan reconstruction in Section IV. The proposed SLAM system utilizes point and plane features, as well as planar regularities, to estimate the camera pose. We refer the reader to [10] for more details on the pose estimation. Next, we focus on describing the operations related to plane landmark management and multi-session map merging.
@@ V-B Localization Accuracy Evaluation
Accurate poses allow planes extracted from different frames to be seamlessly integrated into the global map, maintaining the global consistency of the incrementally reconstructed floorplan. We evaluate the localization performance of our method in large-scale indoor scenes and compare it with the most relevant state-of-the-art visual SLAM systems, including ORB-SLAM3 (stereo mode) [19], Stereo-Plane-SLAM [11], and RSS [10]. To assess the impact of plane features, we also conduct an ablation study using a point-only variant of our system. Experiments are performed on the public VECtor dataset [20] and a self-collected dataset, both containing sequences exceeding 100 meters. The self-collected dataset was captured using a ZED2 camera with GeoSLAM as ground truth (Fig. 5). We align the estimated trajectory with ground truth using the Umeyama algorithm without scaling and use the RMSE of the absolute trajectory error (ATE) to measure global drift. Each sequence is run ten times to mitigate multi-thread randomness, and the median results are reported.
@@
Figure 6: Floorplan reconstruction results on the VECtor Dataset of RoomFormer, VecIM, and Ours. We compute and visualize the mean and RMS Hausdorff error (m) between the reconstructed models and the LiDAR point cloud. The error distribution is color-coded, where blue indicates smaller errors and red represents larger errors.
@@ V-C1 Evaluation on the VECtor Dataset
Figure 7: Floorplan reconstruction results on the self-collected dataset of VecIM, RoomFormer and Ours. For each reconstruction result, the mean Hausdorff error(m) and the error model are computed from the output to the ground truth. From top to bottom, we display the office, café, and meeting rooms, respectively.
@@ V-C2 Evaluation on the Self-collected Dataset
To evaluate the floorplan reconstruction performance of our method in complex indoor environments containing numerous non-lifelong structures (e.g., tables, chairs), and to assess its long-term floorplan reconstruction capability, we used a ZED2 camera to collect data in three large-scale scenes, covering meeting rooms, offices, and a café, among other places, each covering an area of 1000 square meters. In the conference room scenario, due to the inconsistent accessibility of different meeting rooms, we conducted five separate data collections over two days, with each sub-session covering a different region of the scene. The qualitative and quantitative results of floorplan reconstruction on the self-collected dataset are presented in Fig. 7. Our method still reconstructs the most regular and faithful floorplan. Benefiting from the rich textures, the MVS point cloud exhibits relatively low error, enabling VecIM to accurately reconstruct the main structures of the scene. However, due to the presence of non-lifelong objects, VecIM is unable to focus solely on extracting lifelong structures and detects numerous temporary structures such as table sides. Consequently, the reconstructed floorplan contains many redundant elements and still fails to effectively handle navigable areas. Although our method may also extract planes belonging to non-life-long objects during the plane extraction stage, these planes typically do not persist across multiple frames. Therefore, they are not considered valid plane landmarks for the vectorized reconstruction module, which enables our approach to specifically reconstruct life-long structures. Furthermore, the trajectory information is fully leveraged to handle openings in navigable areas, such as corridors and room entrances, ensuring that the reconstructed floorplan accurately represents passable regions. In addition, although our self-collected dataset contains room scenes, RoomFormer lacks generalizability for large-scale environments and thus can only generate a limited number of correctly identified rooms. Notably, by sequentially merging multiple subsequences, our method yields a complete large-scale floorplan that encompasses multiple meeting rooms, exhibiting remarkable accuracy and regularity, which strongly demonstrates the long-term globally consistent mapping capability of our approach. As a result, there is no need to acquire the complete scene data in a single pass, greatly reducing the complexity of data collection and reconstruction.
@@ V-D Runtime Analysis
TABLE IV: Runtime of Each Module in the Proposed Plane Extraction Algorithm
Sequence SFE1 OFE1 MG1 DC1 PPE1 Total
school 20.68 28.96 0.35 0.37 0.24 29.92
Meeting 9.43 13.51 0.19 0.18 0.13 14.01
  • 1
    SFE, OFE, MG, DC, and PPE are the abbreviations for Sobel Feature Extraction, ORB Feature Extraction, Mesh Generation, DBSCAN Clustering, and Plane Parameter Estimation, respectively.
We evaluated the runtime performance of our method on the school sequence from the VECtor dataset and on the Meeting Rooms sequence from our self-collected dataset, comparing the results with ORB-SLAM3. As shown in Table III, we present the runtime of the two threads where our method and ORB-SLAM3 exhibit the most significant differences, while the execution times for the remaining threads are comparable. We utilize parallel CPU threads to independently extract Sobel features and ORB features. Since the former consumes less time, the total plane extraction time is computed as the sum of the ORB feature extraction time and the subsequent plane extraction operations. As shown in Table IV, the runtime of each plane extraction module indicates that ORB feature extraction is the primary time-consuming module. Consequently, our overall plane extraction time is close to that of ORB feature extraction. The floorplan reconstruction module runs on a dedicated thread that operates in parallel with the tracking, local mapping, and loop & map merging threads. It comprises two primary components: segment generation and segment selection. We evaluate the runtime by measuring the duration of the final reconstruction in each sequence, which yields the largest number of segments to be selected and thus represents the most time-consuming execution. Notably, the module leverages plane points provided by valid plane landmarks for floorplan reconstruction, which are typically updated every several frames. As a result, we perform floorplan reconstruction every five frames. This update frequency is sufficient to provide timely updates to the floorplan, thereby enabling real-time floorplan reconstruction.

### v2

@@ I INTRODUCTION
Floorplans are crucial for indoor robot navigation and high-level scene understanding, and are widely applied in scenarios ranging from service robots in assisted living environments to autonomous drones operating in complex indoor facilities. In these scenarios, generating a floorplan in real-time is particularly advantageous, as it not only provides geometric structure for accurate localization but also furnishes semantic cues for higher-level decision-making. Despite the growing need for real-time floorplan reconstruction, existing approaches often present significant limitations. Some methods operate in an offline manner, requiring a complete map of the environment prior to reconstruction [1, 2, 3]. Others rely on expensive sensors such as LiDAR or RGB-D cameras [4, 5], or computationally intensive neural networks [6, 7, 8], which increase both deployment costs and system complexity. While monocular or stereo camera-based SLAM has attempted to use high-level features such as planes for real-time vectorized map construction [9, 10, 11], these methods lack high-level scene layout understanding, thereby failing to achieve complete and compact indoor vectorized reconstruction. Additionally, they cannot perform robust plane extraction in weakly textured scenes.
Figure 1: Floorplan reconstruction results on the meeting rooms sequence from the self-collected dataset by Floorplan-SLAM. Floorplan-SLAM can robustly process stereo images containing low-texture and cluttered objects at 45 FPS on a CPU, and it merges multiple sub-sessions recorded at different times, ultimately achieving real-time, high-accuracy, and long-term floorplan reconstruction.
@@
To address the above problems, an effective real-time solution that works with more accessible and cost-effective sensors, such as stereo cameras, remains uncertain. In this work, we present Floorplan-SLAM, a novel framework that tightly integrates floorplan reconstruction into a multi-session SLAM system by seamlessly interacting with plane extraction, pose estimation, back-end optimization, and loop & map merging. Unlike prior methods, our approach operates exclusively with stereo cameras under standard computing resources, allowing real-time performance in large-scale indoor settings. A robust plane extraction algorithm is first developed to handle weakly textured surfaces by leveraging two spatially complementary features in a compact plane parameter space. We further introduce a floorplan reconstruction module tightly integrated with the SLAM system that formulates and solves an innovative optimization problem using the continuously optimized plane landmarks and poses from the SLAM system, enabling real-time incremental floorplan reconstruction. Furthermore, by leveraging the map merging capability of multi-session SLAM, our method supports long-term floorplan reconstruction across multiple sessions without redundant data collection, making it more efficient for large-scale environments. To the best of the authors’ knowledge, this is the first work that relies solely on a stereo camera and standard CPUs to achieve real-time floorplan reconstruction in large-scale indoor environments, making it well-suited for various robotics applications with strict cost and real-time requirements.
@@
    We introduce a floorplan reconstruction module that is tightly coupled with the multi-session SLAM system, leveraging continuously optimized plane landmarks and poses from the SLAM system to formulate and solve a novel binary linear programming problem with trajectory constraint thereby enabling real-time, high-accuracy, and long-term floorplan reconstruction.
@@ II Related Work
With the development of neural networks, deep learning-based floorplan reconstruction methods have become of interest. These methods [6, 7, 8] typically first project the input dense point cloud onto a top view to create a 2D density map, and then use neural networks to infer the floorplan of the scene from the projected density map. However, these methods are only applicable to specific scenarios, exhibit limited generalizability, and incur high time cost and computational resource consumption.
To achieve real-time floorplan reconstruction, some works [9, 4, 5] utilized visual SLAM to extract plane features from image sequences, which are then used to create landmarks for map construction. However, these methods either lack high-level scene layout understanding [9], thereby failing to achieve complete and compact indoor vectorized reconstruction, or integrate the acquired poses and scene structures in a loosely coupled manner [4, 5], leading to poor system robustness due to the inability to perform mutual optimization between pose and scene structure. In addition, these methods can only process one set of consecutive images at a time, making it impossible to merge multiple sequences and achieve long-term mapping.
@@
Figure 2: System Overview. Floorplan-SLAM tightly integrates a stereo point-plane SLAM system with a floorplan reconstruction module. The former provides accurate plane landmarks and trajectory information, while the latter formulates and solves a novel optimization problem, ultimately achieving real-time, high-accuracy, and long-term floorplan reconstruction.
@@ III Point-Plane-Based Stereo SLAM System
In this section, we introduce the proposed point-plane-based stereo SLAM system. The system first uses two sets of spatially complementary point features to extract planes, and then employs both point and plane features to estimate poses. Plane landmarks are subsequently created and managed for floorplan reconstruction. In addition, a multi-session map merging mechanism is introduced to support long-term mapping. The overview of the system is illustrated on the left side of Fig. 2. We refer the reader to [10] for more details on the pose estimation. Next, we focus on describing the operations related to plane extraction, plane landmark management, and multi-session map merging.
@@ IV Floorplan Reconstruction
The floorplan reconstruction module runs in parallel with the tracking, local mapping, and loop & map merging threads. It processes the plane landmarks and trajectory information generated by the SLAM system and reconstructs the floorplan by formulating and solving a novel optimization problem. An overview of this workflow is shown on the right side of Fig. 2.
@@ V-B Localization Accuracy Evaluation
Accurate poses allow planes extracted from different frames to be seamlessly integrated into the global map, maintaining the global consistency of the incrementally reconstructed floorplan. We evaluate the localization performance of our method in large-scale indoor scenes and compare it with the most relevant state-of-the-art visual SLAM systems, including ORB-SLAM3 (stereo mode) [19], Stereo-Plane-SLAM [11], and RSS [10]. To assess the impact of plane features, we also conduct an ablation study using a point-only variant of our system. Experiments are performed on the public VECtor dataset [20] and a self-collected dataset, both containing sequences exceeding 100 meters. The self-collected dataset was captured using a ZED2 camera with GeoSLAM ZEB as ground truth (Fig. 5). We align the estimated trajectory with ground truth using the Umeyama algorithm without scaling and use the RMSE of the absolute trajectory error (ATE) to measure global drift. Each sequence is run ten times to mitigate multi-thread randomness, and the median results are reported.
@@
Figure 6: Floorplan reconstruction results on the VECtor dataset of RoomFormer, VecIM, and Ours. We compute and visualize the mean and RMS Hausdorff error (m) between the reconstructed models and the LiDAR point cloud ground truth. The error distribution is color-coded, where blue indicates smaller errors and red represents larger errors.
@@ V-C1 Evaluation on the VECtor Dataset
Figure 7: Floorplan reconstruction results on the self-collected dataset of VecIM, RoomFormer and Ours. For each reconstruction result, the mean Hausdorff error (m) and the error model are computed from the output to the ground truth. From top to bottom, we display the office, café, and meeting rooms, respectively. Our method achieves the best performance in both reconstruction accuracy and regularity.
@@ V-C2 Evaluation on the Self-collected Dataset
To evaluate the floorplan reconstruction performance of our method in complex indoor environments containing numerous non-lifelong structures (e.g., tables, chairs), and to assess its long-term floorplan reconstruction capability, we used a ZED2 camera to collect data in three large-scale scenes, covering meeting rooms, offices, and a café, among other places, each covering an area of 1000 square meters. In the meeting rooms scenario, due to the inconsistent accessibility of different conference rooms, we conducted five separate data collections over two days, with each sub-session covering a different region of the scene, as shown in Fig. 1. The qualitative and quantitative results of floorplan reconstruction on the self-collected dataset are presented in Fig. 7. Our method still reconstructs the most regular and faithful floorplan. Benefiting from rich textures, the MVS point cloud exhibits relatively low error, allowing VecIM to accurately reconstruct main structures. However, VecIM is unable to focus solely on extracting lifelong structures and detects numerous temporary structures such as table sides, adding redundant elements and failing to handle navigable areas effectively. In contrast, although our method may initially detect planes from non-lifelong objects, these planes typically do not persist across multiple frames and are not considered valid plane landmarks for the floorplan reconstruction module, enabling a focus on lifelong structures. Moreover, by fully leveraging trajectory information, we effectively handle openings in navigable areas, such as corridors and room entrances, ensuring that the reconstructed floorplan accurately represents passable regions. In addition, although our self-collected dataset contains room scenes, RoomFormer lacks generalizability for large-scale environments and thus can only generate a limited number of correctly identified rooms. Notably, by sequentially merging multiple subsequences, our method yields a complete large-scale floorplan that encompasses multiple meeting rooms, exhibiting remarkable accuracy and regularity, which strongly demonstrates the long-term globally consistent mapping capability of our approach. As a result, there is no need to acquire the complete scene data in a single pass, greatly reducing the complexity of data collection and reconstruction.
@@ V-D Runtime Analysis
The runtime analysis was conducted on a desktop with an AMD Ryzen 7 4800H CPU (running at 2.9 GHz) and 32 GB of RAM. We evaluated our method on the school sequence from the VECtor dataset and on the meeting rooms sequence from our self-collected dataset, comparing the results with ORB-SLAM3. Table III shows the runtimes of the two threads where our method and ORB-SLAM3 differ most; the other threads show comparable execution times. We extract Sobel and ORB features in parallel CPU threads, with the former being faster. Moreover, ORB extraction is the main time-consuming step compared to other plane extraction operations, such as mesh generation, clustering, and plane parameter estimation. As a result, our overall plane extraction time is close to that of ORB feature extraction. The floorplan reconstruction module comprises two primary components: segment generation and selection. We evaluate the runtime by measuring the duration of the final reconstruction in each sequence, which is the most time-consuming due to the largest number of segments. Notably, considering that valid plane landmarks are typically updated every several frames, we perform floorplan reconstruction every five frames. This frequency is sufficient to provide timely updates, thereby enabling real-time floorplan reconstruction. We also recorded the total reconstruction time consumed by VecIM, including the MVS dense point cloud reconstruction and the floorplan reconstruction stages. The first stage took 16 hours and 38.9 minutes, and the second stage took 5.1 minutes, resulting in a total of 16 hours and 44 minutes. In contrast, we processed only stereo images at 45 FPS, and completed the entire floorplan reconstruction in 9.4 minutes, which is two orders of magnitude faster than VecIM. This result strongly demonstrates the efficiency and practicality of our method.